How many hydrogen atoms are in 25.0 g of water?


1
Expert's answer
2021-03-03T06:30:36-0500

number of moles of water = 25g/16gmol-¹

= 1.563mol


H2O(l) -> 2H+(g) + ½O2-(g)

Therefore in 1.563 moles of water, there are 3.125 moles of hydrogen atom


3.125 moles of hydrogen atom = 3.125 (6.02×10²³) = 1.88 × 10²⁴ hydrogen atoms.